1. Matt loses his job, but because he starts to receive unemployment insurance, he initially decides to limit his job search to those jobs that are most specific to his interest in sailing around the world. Other things the same, the unemployment rate

   increases, and the labor-force participation rate is unaffected.

   increases, and the labor-force participation rate decreases.

   decreases, and the labor-force participation rate increases.

   decreases, and the labor-force participation rate decreases.

2. Use the balance sheet for the following questions.

Table 29-3

Last Bank of Cedar Bend

Assets Liabilities
Required Reserves $7,500 Deposits $150,000
Loans $142,500



Refer to Table 29-3. The reserve ratio is

   1 percent.

   5 percent.

   10 percent.

   20 percent.

3. Which of the following lists two things that both increase the money supply?

   raise the discount rate, make open market purchases

   raise the discount rate, make open market sales

   lower the discount rate, make open market purchases

   lower the discount rate, make open market sales
Answer #1

1. increases, and the labor-force participation rate is unaffected.

Explanation: Matt becomes unemployed but he stays in the workforce.

2. 5 percent.

Explanation: Reserve ratio = required reserve/ deposit * 100 = 7,500/150,000 * 100 = 1/20 * 100 = 5%

3. lower the discount rate, make open market purchases

Explanation: When the discount rate is lowered, the money supply increases. Also, when the Fed makes open market purchases, money flows from the Fed to the banks.